how many liters of a 5% acidic solution must be mixed with a 20% acidic solution to make 100 L having a 14% concentration
Answer by stanbon(75887)

how many liters of a 5% acidic solution must be mixed with a 20% acidic solution to make 100 L having a 14% concentration
------
acid + acid = acid
0.05x + 0.20(100-x) = 0.14*100
------------------------------------
5x + 20*100 - 20x = 14*100
-------
-15x = -6*100
-------
x = 40 L (amt. of 5% solution needed)
100-x = 60 L (amt. of 20% solution needed)
-----------
Cheers,
Stan H.
-----------